OpenStudy (iloveleatiforever):

Solve for k. 3k – 5 = 4

OpenStudy (tuxadog):

k=2

OpenStudy (jailbird):

3(2)-5 does not = 4

OpenStudy (iloveleatiforever):

ok

OpenStudy (tuxadog):

oops I typed the wrong number sorry one sec

OpenStudy (jailbird):

3(3)-5=4

OpenStudy (jailbird):

3(3)=9 9-5=4

OpenStudy (tuxadog):

k=3